OpenStudy (anonymous):

Evaluate: cos(Arctan 15/8)

OpenStudy (nikita2):

tan = sin/cos. sin(arctan15/8)/cos(arctan15/8) = 15/8 and sin^2(arctan15/8) + cos^2(arctan15/8) =1. So you have two equations. try to solve them)

OpenStudy (anonymous):

i appreciate the help but i still cannot figure this out.

myininaya (myininaya):

hey you still there?

OpenStudy (nikita2):

let's arctan15/8 = a. We have sin(a) = 15/8cos(a) and cos^2(a) = sin^2(a). Then cos^2(a) = 1 - 225/64cos^2(a). So cos^2(a) = 1/(1 - 225/64) or \[\cos(a) = \sqrt{1/(1 + 225/64 )}\]

myininaya (myininaya):

i kept writing 18 for some reason so ignore the scratch outs

myininaya (myininaya):

you can draw a right traingle they are pretty easy to work with and label the sides based on what you have up there

OpenStudy (anonymous):

alright now it makes sense, thanks a lot guys